OpenStudy (anonymous):

Look at the equations. If 3a + 2(b − 2c) = p then p = 3a + 2(b − 2c) Which property is used to write the equations? Answer associative property of addition commutative property of addition distributive property of equality reflexive property of equality

Parth (parthkohli):

Reflexive property. Do you know about the other three?

OpenStudy (anonymous):

That's what I thought

OpenStudy (anonymous):

But I doubted myself

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow Associative - (a+b)+c = a + (b + c) }\) \(\Large \color{purple}{\rightarrow Distributive - a(b + c) = ab + ac }\) \(\Large \color{purple}{\rightarrow Commutative - a + b = b + a }\)

OpenStudy (apoorvk):

Reflexive property says: if all toogs are donks, then all donks are toogs as well. ;) Basically, if x=y, then y=x

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow Reflexive - if a = b, b = a }\)

Parth (parthkohli):

@apoorvk lol I thought you'd gone for a 10-day stroll in the park

OpenStudy (apoorvk):

The 'park' was so irritating, I thought this 'wilderness' is better (for this evening atleast) ;)

OpenStudy (anonymous):

Thanks guys!

Parth (parthkohli):

Hahaha @apoorvk @evergirl yw :)